A fire engine is being driven away from you at a speed of 15.4 m/s. One of the notes in its siren sequence has a fundamental frequency of 244 Hz. If the speed of sound is 338 m/s, the fundamental frequency of that particular note will seem to you to be ___Hz. (Give your answer with the correct number of sig digs and do not include units)



1
Expert's answer
2021-07-21T09:05:43-0400

Given:

"v_s=15.4\:\rm m/s"

"f_0=244\:\rm Hz"

"u=338\:\rm m/s"


"f=f_0\frac{u}{u+v_s}=244*\frac{338}{338+15.4}=214\:\rm Hz"